# UNIFY: Unified Index for Range Filtered Approximate Nearest Neighbors Search

### v1

@@ Abstract.
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
@@ 1. Introduction
(5) Range-aware search strategy selection (O4). Inspired by the effectiveness of Strategies A, B, and C for different query ranges, we developed a heuristic for range-aware strategy selection. Let $Y$ denote the cardinality of objects that fall within the query range, our heuristic is: use Strategy A for $Y\leq\tau_{A}$, Strategy B for $Y\geq\tau_{B}$, and Strategy C for $\tau_{A}<Y<\tau_{B}$. Here $\tau_{A}$ and $\tau_{B}$ serve as thresholds to distinguish the ranges for which each strategy is most effective, and they can be derived from history data statistics. In this paper, we run a set of sample queries to collect statistics and determine $\tau_{A}$ and $\tau_{B}$. Experimental results show that this heuristic is effective.
@@ 2 $\mathbb{G}[v]\leftarrow$ Prune($v$, $ann$, $M$)
Hierarchical Navigable Small World Graph (HNSW) (Malkov and Yashunin, 2018). HNSW is inspired by the 1D probabilistic structure of the skip list (Pugh, 1990), where each layer is a linked list ordered by 1D values. The bottom layer includes all data objects, while the upper layers contain progressively fewer objects. HNSW extends this structure by replacing linked lists with PGs, enabling efficient hierarchical search. The hierarchical structure of HNSW is similar to that in Figure 4. As outlined in Algorithm 1, HNSW uses a greedy approach to find the $k$NN for query vector $q$ in each layer. The neighbors of each node $v_{i}$ in a given layer $\mathbb{G}$ are stored in adjacency lists $\mathbb{G}[v_{i}]$. The search starts at an entry point $ep$, the most recently inserted vector at the topmost layer (Malkov and Yashunin, 2018). It repeatedly selects the nearest object $o$ to $q$ from $cand$ (Line 5), adds $o$’s unvisited neighbors to $cand$ (Lines 8–12), and updates $ann$ (Lines 11–13). The search terminates when all nodes in $cand$ are farther from $q$ than those in $ann$. As shown in Algorithm 2, the hierarchical search in HNSW begins at a coarse layer to identify promising regions and progressively descends to finer layers for detailed exploration. To improve accuracy, Algorithm 2 uses the parameter $ef$ ($ef>k$), initially exploring $ef$ neighbors to broaden the search region, and finally returns the top-$k$ results.
HNSW is built incrementally based on Algorithm 3. When inserting a new object $v$ into an HNSW layer, the process begins by searching for its top $efCons$ nearest neighbors using the ANNSearch algorithm (Line 1). The Prune heuristic is then applied to limit $v$’s connections to a maximum of $M$ (Line 2). The value of $M$, typically set between 5 and 48 (Malkov and Yashunin, 2018), balances accuracy and efficiency, with its optimal value determined experimentally. The Prune method initially sorts $v$’s candidate neighbors by distance. For each candidate $r$, if there exists a neighbor $e$ that satisfies $\Gamma(v,r)>\Gamma(e,r)$, $r$ is pruned. Otherwise, an edge between $v$ and $r$ is inserted. The process continues until $v$ has $M$ neighbors. Additionally, $v$ is added to the neighbor lists of its identified neighbors, with the Prune method executed for each to maintain the $M$ connection limit (Lines 3–5). The Prune strategy prevents the graph from becoming overly dense, ensuring efficient navigation. HNSW achieves state-of-the-art ANNS time complexity of $O(\log n)$ (Wang et al., 2021a) and demonstrates top-tier practical performance indicated by various benchmarks (Li et al., 2019; Aumüller et al., 2020; Wang et al., 2021a). It also serves as the backbone for various vector databases such as Pinecone (Pinecone, 2021), Weaviate (Weaviate, 2019), and Milvus (Wang et al., 2021b).
@@ 3. Segmented Inclusive Graph
We aim to develop a PG-based index that integrates three search strategies and supports incremental construction. We start by discussing the design of a graph optimized for hybrid filtering. In Section 3.1, we introduce the Segmented Inclusive Graph (SIG), a novel graph family that leverages attribute segmentation and graph inclusivity for efficient hybrid filtering. Section 3.2 discusses the challenges of constructing an SIG using the basic PG structure $k$NNG as a case study and presents our practical solutions.
@@ 3.1. SIG Overview
Attribute Segmentation. Given a dataset of $n$ objects, RF-ANNS is concerned only with the $n^{\prime}$ objects within the query range, where $n^{\prime}\leq n$. To efficiently filter out the qualified objects, we employ the attribute segmentation method to reduce the search space. Assuming that the attribute distribution remains stable, we first sample objects from the dataset and sort them by their attribute values, then apply an equi-depth histogram (Piatetsky-Shapiro and Connell, 1984) to partition them. The histogram bin boundaries define the attribute intervals for each segment, partitioning the dataset into disjoint subsets with similar size for further indexing.
@@
Based on Definition 1, we can derive an exhaustive SIG construction algorithm: build PGs for all possible segment combinations and merge them into a single graph. While this guarantees graph inclusivity, it faces significant computational challenges. With $S$ segments, the number of combinations is $\sum_{i=1}^{S}\binom{S}{i}=2^{S}-1$, and the space complexity of SIG scales as $O(2^{S}-1)$, making the construction of all PGs both computationally and spatially prohibitive [^1] In fact, for RF-ANNS, we only need to consider continuous segments, with potential combinations totaling $0.5(n^{2}-n)$, which is also prohibitive. Since the conclusions in this paper apply to both all segment combinations and continuous ones, we discuss the harder scenario, i.e., all segment combinations, to maintain generality. . Thus, we propose a space-efficient SIG index that scales as $O(S)$ in the next section.
@@ 3.2. SIG-$k$NNG
In this section, we conduct a case study on the basic PG structure $k$NNG to demonstrate the limitations of the exhaustive method. Then, we introduce SIG-$k$NNG, a novel graph structure that guarantees inclusivity without exhaustively building all possible $k$NNGs. Limitation of the Exhaustive Method. For each object $v$ in a dataset $\mathcal{D}$, constructing a $k$NNG involves adding a directed edge $(v,o)$ for each object $o\in kNN(v,\mathcal{D}\setminus\{v\})$. When the dataset is divided into $S$ subsets by attribute segmentation, the exhaustive method requires running the construction algorithm $12^{S}-$ 1 times. Figure 1 illustrates an example. Here, the dataset $\mathcal{D}$ is divided into two subsets $\mathcal{D}_{1}=\{v_{1},v_{2},v_{3},v_{4}\}$ and $\mathcal{D}_{2}=\{v_{5},v_{6},v_{7},v_{8}\}$. We need to build $k$NNGs ($k$=3) for $\mathcal{D}_{1}$, $\mathcal{D}_{2}$, and $\mathcal{D}_{1}\cup\mathcal{D}_{2}$, respectively. Such exhaustive construction is unnecessary, as the $k$NN of a union is inherently within the individual $k$NN sets. For example, all objects in $kNN(v_{4},\mathcal{D}_{1}\cup\mathcal{D}_{2})$ can be found in $kNN(v_{4},\mathcal{D}_{1})$cup $kNN(v_{4},\mathcal{D}_{2})$. In other words, $v_{4}$’s three edges in graph ③ are already included in graphs ① and ②, making the construction of graph ③ redundant.
Structure of SIG-$k$NNG. Inspired by the prior observations, we introduce SIG-$k$NNG, a novel graph structure that achieves inclusivity without the need for exhaustive $k$NNG construction. SIG-$k$NNG uses a segmented adjacency list to storage the outgoing edges for each object in the graph. As illustrated in Figure 2, given a dataset $\mathcal{D}$ segmented into $S$ subsets $\mathcal{D}_{1},\dots,\mathcal{D}_{S}$, the adjacency list for any object $v$ is divided into $S$ chunks. For example, the full adjacency list $\mathbb{G}[v_{4}]$ of object $v_{4}$ is segmented into two chunks $\mathbb{G}[v_{4}][1]$ and $\mathbb{G}[v_{4}][2]$. The $i$-th chunk stores only $v$’s $k$NN within $\mathcal{D}_{i}$. Based on this design, to add SIG-$k$NNG edges for $v$, we perform $k$NN searches $S$ times to find its neighbors in each subset, instead of searching $12^{S}-$ 1 times across all subset combinations. In the following, we introduce the formal definition of SIG-$k$NNG and prove that SIG-$k$NNG can exactly guarantee inclusivity.
@@
Based on Lemma 4, we derive the inclusivity of SIG-$k$NNG, as shown in Theorem 5. The proof is omitted, as it is straightforward to follow.
@@ 4.1. HSIG Construction
Backbone Graph for Hybrid Filtering. Based on Theorem 5, we present HSIG’s backbone graph, which is designed to ensure inclusivity regarding HNSW and support efficient hybrid filtering. As depicted in Algorithm 3, when inserting a new object $v$ into an HNSW layer, the core operation involves finding $v$’s $efCons$ nearest neighbors and establishing up to $M$ connections. This procedure is similar to $k$NNG’s construction, which finds $k$NN and establish at most $k$ connections. Therefore, we use the segmented adjacency list introduced by SIG-$k$NNG to guarantee the inclusivity for all possible HNSW connections. With $S$ segments, connections are stored in $S$ chunks, with a maximum degree of $M$ per chunk. The values of $S$ and $M$ are experimentally determined, with default values set to 8 and 16, respectively. Since each chunk contains the HNSW connections in the corresponding segment, the objects and connections in chunk $j$ form a sub-graph $\mathbb{G}_{j}$ of the backbone graph in HSIG. When inserting object $v$ that is located in segment $i$, Algorithm 4 outlines how to build $v$’s connections in segment $j$ at an HSIG layer. First, we search for $v$’s $efCons$ nearest neighbors in $\mathbb{G}_{j}$ (Line 1) using ANNSearch (Algorithm 1), then select up to $M$ neighbors using the Prune algorithm. For each neighbor $o$, we establish mutual connections between $v$ and $o$ within their respective segments (Lines 2–4). If the number of neighbors of $o$ in segment $i$ exceeds $M$, we apply the Prune algorithm to discard the extra neighbors (Lines 5–6). This algorithm guarantees the approximate inclusivity of HSIG, as demonstrated in Section 5.4.
@@ 9                   $\mathcal{N}^{\prime}\leftarrow$ Prune($o$, $\mathbb{G}[o]$, $M$);
Next, we present the complete insertion algorithm to construct an HSIG. The structure of HSIG can be seen as a set of HNSWs for different segments, augmented with auxiliary structures for pre- and post-filtering. The construction consists of three main steps: building the HNSW in each segment, adding skip list connections, and achieving global edge masking. Each step is performed hierarchically and incrementally, as outlined in Algorithm 6. First, we find the segment to which object $v$ belongs (Line 1). The maximum layer $level$ of $v$ is determined randomly using an exponentially decaying probability distribution normalized by $m_{L}$ (Line 2; see (Malkov and Yashunin, 2018) for more details). We traverse all segments (Line 3) to sequentially establish the HNSW $\mathbb{HG}_{j}$ in each segment $j$. We determine the maximum level $L$ and the entry point $ep$ of $\mathbb{HG}_{j}$ (Lines 4–5). For each layer $l$ from $L$ to $level+1$, we use ANNSearch (Algorithm 1) to find the nearest neighbor $ep$ of $v$ in $\mathbb{HG}_{j}^{l}$, where $\mathbb{HG}_{j}^{l}$ denotes layer $l$ in $\mathbb{HG}_{j}$ (Lines 6–7). For each layer $l$ from $level$ to 0, we employ BackboneConnectionsBuild (Algorithm 4) to insert $v$ into $\mathbb{HG}_{j}^{l}$ and use the nearest neighbor of $v$ as the entry point for the next layer (Lines 8–10). After building $\mathbb{HG}_{j}$, we update its entry point if necessary (Line 11) based on the rule that the first object in the topmost layer becomes the entry point. Next, $v$ is inserted into the skip list using the method described in (Pugh, 1990) (Line 12). Finally, we use GlobalEdgeMasking (Algorithm 5) to mask unnecessary edges for post-filtering from layer 0 to $level$ (Lines 13–14).
@@ 4.2. Search on HSIG
Strategy B (Post-Filtering). We employ the hierarchical search scheme outlined in Algorithm 2 to implement ANNS and then filter out objects that fall within the query range. The ANNS is performed over a global HNSW of the entire dataset, with edges marked by bitmaps. During ANNS, we only consider outgoing edges marked as 1 in the object’s bitmap. The search retrieves the top-$ef$ nearest neighbors ($ef>k$) in the bottom layer. We then perform attribute filtering on these top-$ef$ results to obtain the final top-$k$ results.
@@
Range-aware Strategy Selection. Let $Y$ be the cardinality of objects within a query range. Observing that pre-, post-, and hybrid filtering outperform for small-, large-, and mid-range queries, respectively, we propose the following heuristic for strategy selection: use Strategy A if $Y\leq\tau_{A}$, Strategy B if $Y\geq\tau_{B}$, and Strategy C if $\tau_{A}<Y<\tau_{B}$. Here, $\tau_{A}$ and $\tau_{B}$ are thresholds that distinguish the optimal ranges for each strategy and can be derived from statistical analysis of historical data. Given these thresholds, we can estimate the cardinality of an incoming query to apply the heuristic. Since both statistic collection and cardinality estimation are well studied (Harmouch and Naumann, 2017; Oommen and Rueda, 2002) and are not the focus of this paper, we provide a simple preprocessing method to validate the effectiveness of our heuristic. Specifically, we sample a set of objects from the base dataset to use as queries and assign each a random query range. We then run these queries using the three strategies and record their recall and latency metrics. Given a recall target, we analyze records that meet this requirement, identifying two turning points where pre- and post-filtering outperform hybrid filtering. These two points establish $\tau_{A}$ and $\tau_{B}$, guiding strategy selection for future queries.

### v2

@@ ABSTRACT.
PVLDB, 18(4): 1118 - 1130, 2024.
doi:10.14778/3717755.3717770 This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 18, No. 4 ISSN 2150-8097.
doi:10.14778/3717755.3717770
@@ 1. INTRODUCTION
(5) Range-aware search strategy selection (O4). Inspired by the effectiveness of Strategies A, B, and C for different query ranges, we developed a heuristic for range-aware strategy selection. Let $Y$ denote the cardinality of objects that fall within the query range, our heuristic is: use Strategy A for $Y\leq\tau_{A}$, Strategy B for $Y\geq\tau_{B}$, and Strategy C for $\tau_{A}<Y<\tau_{B}$. Here, $\tau_{A}$ and $\tau_{B}$ serve as thresholds to distinguish the ranges for which each strategy is most effective, and they can be derived from historical data statistics. In this paper, we run a set of sample queries to collect statistics and determine $\tau_{A}$ and $\tau_{B}$. Experimental results show that this heuristic is effective.
@@ 2 $\mathbb{G}[v]\leftarrow$ Prune($v$, $ann$, $M$)
Hierarchical Navigable Small World Graph (HNSW) (Malkov and Yashunin, 2018). HNSW is inspired by the 1D probabilistic structure of the skip list (Pugh, 1990), where each layer is a linked list ordered by 1D values. The bottom layer includes all objects, while the upper layers contain progressively fewer objects. HNSW extends this structure by replacing linked lists with PGs, enabling efficient hierarchical search. The hierarchical structure of HNSW is similar to that in Figure 4. As shown in Algorithm 1, HNSW uses a greedy approach to find the $k$NN for query vector $q$ in each layer. The neighbors of each node $v_{i}$ in a given layer $\mathbb{G}$ are stored in adjacency lists $\mathbb{G}[v_{i}]$. The search starts at an entry point $ep$, the most recently inserted vector at the topmost layer (Malkov and Yashunin, 2018). It repeatedly selects the nearest object $o$ to $q$ from $cand$ (Line 5), adds $o$’s unvisited neighbors to $cand$ (Lines 8–12), and updates $ann$ (Lines 11–13). The search ends when all nodes in $cand$ are farther from $q$ than those in $ann$. As shown in Algorithm 2, the hierarchical search in HNSW begins at a coarse layer to identify promising regions and progressively descends to finer layers for detailed exploration. To improve accuracy, Algorithm 2 uses the parameter $ef$ ($ef>k$), initially exploring $ef$ neighbors to broaden the search region, and finally returns the top-$k$ results.
HNSW is built incrementally based on Algorithm 3. When inserting a new object $v$ into an HNSW layer, the process begins by searching for its top $efCons$ nearest neighbors using the ANNSearch algorithm (Line 1). The Prune heuristic is then applied to limit $v$’s connections to a maximum of $M$ (Line 2). The value of $M$, typically set between 5 and 48 (Malkov and Yashunin, 2018), balances accuracy and efficiency, with its optimal value determined experimentally. The Prune method initially sorts $v$’s candidate neighbors by distance. For each candidate $r$, if a neighbor $e$ satisfies $\Gamma(v,r)>\Gamma(e,r)$, $r$ is pruned. Otherwise, an edge between $v$ and $r$ is inserted. The process continues until $v$ has $M$ neighbors. Additionally, $v$ is added to the neighbor lists of its identified neighbors, and the Prune method is applied to each neighbor to maintain the $M$-neighbor limit (Lines 3–5). The Prune strategy prevents the graph from becoming overly dense, ensuring efficient navigation. HNSW achieves state-of-the-art ANNS time complexity of $O(\log n)$ (Wang et al., 2021a) and demonstrates top-tier practical performance indicated by various benchmarks (Li et al., 2019; Aumüller et al., 2020; Wang et al., 2021a). It also serves as the backbone for various vector databases such as Pinecone (Pinecone, 2021), Weaviate (Weaviate, 2019), and Milvus (Wang et al., 2021b).
@@ 3. SEGMENTED INCLUSIVE GRAPH
We aim to develop a PG-based index that integrates three search strategies and supports incremental construction. Section 3.1 introduces the Segmented Inclusive Graph (SIG), a novel graph family that leverages attribute segmentation and inclusivity for hybrid filtering. Section 3.2 explores the challenges of constructing an SIG using the basic PG structure $k$NNG and presents practical solutions.
@@ 3.1. SIG Overview
Attribute Segmentation. Given a dataset of $n$ objects, RF-ANNS is concerned only with the $n^{\prime}$ objects within the query range, where $n^{\prime}\leq n$. To efficiently filter out the qualified objects, we employ the attribute segmentation method to reduce the search space. Assuming that the attribute distribution remains stable, we first sample objects from the dataset and sort them by their attribute values, then apply an equi-depth histogram (Piatetsky-Shapiro and Connell, 1984) to partition them. The histogram bin boundaries define the attribute intervals for each segment, partitioning the dataset into disjoint subsets with similar sizes for further indexing.
@@
Based on Definition 1, we derive an exhaustive SIG construction algorithm: build PGs for all segment combinations and merge them into a single graph. While it guarantees graph inclusivity, it faces significant computational challenges. With $S$ segments, the number of combinations is $\sum_{i=1}^{S}\binom{S}{i}=2^{S}-1$, and the space complexity of SIG scales as $O(2^{S}-1)$, making the construction of all PGs both computationally and spatially prohibitive [^1] In fact, for RF-ANNS, we only need to consider continuous segments, with potential combinations totaling $0.5(n^{2}-n)$, which is also prohibitive. Since the conclusions in this paper apply to both all segment combinations and continuous ones, we discuss the harder scenario, i.e., all segment combinations, to maintain generality. . Thus, we propose a space-efficient SIG index that scales as $O(S)$ in the next section.
@@ 3.2. SIG-$k$NNG
In this section, we conduct a case study on $k$NNG to demonstrate the limitations of the exhaustive method. Then, we introduce SIG-$k$NNG, a novel graph structure that guarantees inclusivity without exhaustively building all possible $k$NNGs.
Limitation of the Exhaustive Method. For each object $v$ in a dataset $\mathcal{D}$, constructing a $k$NNG involves adding a directed edge $(v,o)$ for each object $o\in kNN(v,\mathcal{D}\setminus\{v\})$. When the dataset is divided into $S$ subsets by attribute segmentation, the exhaustive method requires running the construction algorithm $12^{S}-$ 1 times. Figure 1 illustrates an example. Here, the dataset $\mathcal{D}$ is divided into two subsets $\mathcal{D}_{1}=\{v_{1},v_{2},v_{3},v_{4}\}$ and $\mathcal{D}_{2}=\{v_{5},v_{6},v_{7},v_{8}\}$. We need to build $k$NNGs ($k$=3) for $\mathcal{D}_{1}$, $\mathcal{D}_{2}$, and $\mathcal{D}_{1}\cup\mathcal{D}_{2}$, respectively. Such exhaustive construction is unnecessary, as the $k$NN of a union is inherently within the individual $k$NN sets. For example, all objects in $kNN(v_{4},\mathcal{D}_{1}\cup\mathcal{D}_{2})$ can be found in $kNN(v_{4},\mathcal{D}_{1})$cup $kNN(v_{4},\mathcal{D}_{2})$. In other words, $v_{4}$’s three edges in graph ③ are already included in graphs ① and ②, making the construction of graph ③ redundant.
Structure of SIG-$k$NNG. Inspired by the prior observations, we introduce SIG-$k$NNG, a novel graph structure that achieves inclusivity without the need for exhaustive $k$NNG construction. SIG-$k$NNG uses a segmented adjacency list to store the outgoing edges for each object in the graph. As illustrated in Figure 2, given a dataset $\mathcal{D}$ segmented into $S$ subsets $\mathcal{D}_{1},\dots,\mathcal{D}_{S}$, the adjacency list for any object $v$ is divided into $S$ chunks. For example, the full adjacency list $\mathbb{G}[v_{4}]$ of object $v_{4}$ is segmented into two chunks $\mathbb{G}[v_{4}][1]$ and $\mathbb{G}[v_{4}][2]$. The $i$-th chunk stores only $v$’s $k$NN within $\mathcal{D}_{i}$. Based on this design, to add SIG-$k$NNG edges for $v$, we perform $k$NN searches $S$ times to find its neighbors in each subset, instead of searching $12^{S}-$ 1 times across all subset combinations. In the following, we introduce the formal definition of SIG-$k$NNG and prove that SIG-$k$NNG can exactly guarantee inclusivity.
@@
Based on Lemma 4, we derive the inclusivity of SIG-$k$NNG, as presented in Theorem 5. The proof is omitted here for brevity, as it is straightforward to follow.
@@ 4.1. HSIG Construction
Backbone Graph for Hybrid Filtering. Based on Theorem 5, we present HSIG’s backbone graph, which is designed to ensure inclusivity regarding HNSW and support efficient hybrid filtering. As depicted in Algorithm 3, when inserting a new object $v$ into an HNSW layer, the core operation involves finding $v$’s $efCons$ nearest neighbors and establishing up to $M$ connections. This procedure is similar to $k$NNG’s construction, which finds $k$NN and establishes at most $k$ connections. Therefore, we use the segmented adjacency list introduced by SIG-$k$NNG to guarantee the inclusivity for all possible HNSW connections. With $S$ segments, connections are stored in $S$ chunks, with a maximum degree of $M$ per chunk. The values of $S$ and $M$ are experimentally determined, with default values set to 8 and 16, respectively. Since each chunk contains the HNSW connections in the corresponding segment, the objects and connections in chunk $j$ form a sub-graph $\mathbb{G}_{j}$ of the backbone graph in HSIG. When inserting object $v$ that is located in segment $i$, Algorithm 4 outlines how to build $v$’s connections in segment $j$ at an HSIG layer. First, we search for $v$’s $efCons$ nearest neighbors in $\mathbb{G}_{j}$ (Line 1) using ANNSearch (Algorithm 1), then select up to $M$ neighbors using the Prune algorithm. For each neighbor $o$, we establish mutual connections between $v$ and $o$ within their respective segments (Lines 2–4). If the number of neighbors of $o$ in segment $i$ exceeds $M$, we apply the Prune algorithm to discard the extra neighbors (Lines 5–6). This algorithm guarantees the approximate inclusivity of HSIG, as demonstrated in Section 5.4.
@@ 9                   $\mathcal{N}^{\prime}\leftarrow$ Prune($o$, $\mathbb{G}[o]$, $M$);
Next, we present the complete insertion algorithm for constructing an HSIG. The HSIG structure consists of a set of HNSWs for different segments, augmented with auxiliary structures for pre- and post-filtering. Construction involves three main steps: building the HNSW in each segment, adding skip list connections, and applying global edge masking. Each step is performed hierarchically and incrementally, as outlined in Algorithm 6. First, we find the segment to which object $v$ belongs (Line 1). The maximum layer $level$ of $v$ is randomly assigned using an exponentially decaying probability distribution normalized by $m_{L}$ (Line 2; see (Malkov and Yashunin, 2018) for details). We then traverse all segments (Line 3) to sequentially build the HNSW $\mathbb{HG}_{j}$ in each segment $j$, determining its maximum level $L$ and entry point $ep$ (Lines 4–5). For each layer $l$ from $L$ to $level+1$, we use ANNSearch (Algorithm 1) to find the nearest neighbor $ep$ of $v$ in $\mathbb{HG}_{j}^{l}$, where $\mathbb{HG}_{j}^{l}$ denotes layer $l$ in $\mathbb{HG}_{j}$ (Lines 6–7). Then, for each layer $l$ from $level$ to 0, we employ BackboneConnectionsBuild (Algorithm 4) to insert $v$ into $\mathbb{HG}_{j}^{l}$, using its nearest neighbor as the entry point for the next layer (Lines 8–10). After building $\mathbb{HG}_{j}$, we update its entry point if necessary (Line 11), ensuring the first object in the topmost layer serves as the entry. Next, $v$ is inserted into the skip list using the method in (Pugh, 1990) (Line 12). Finally, we use GlobalEdgeMasking (Algorithm 5) to mask unnecessary edges for post-filtering from layer 0 to $level$ (Lines 13–14).
@@ 4.2. Search on HSIG
Strategy B (Post-Filtering). We employ the hierarchical search scheme in Algorithm 2 for ANNS, followed by filtering objects within the query range. ANNS is performed over a global HNSW with edges marked by bitmaps. During the search, only outgoing edges marked as 1 in an object’s bitmap are considered. The search retrieves the top-$ef$ nearest neighbors ($ef>k$) in the bottom layer, after which attribute filtering is applied to obtain the top-$k$ results.
@@
Range-aware Strategy Selection. Let $Y$ be the cardinality of objects within a query range. Observing that pre-, post-, and hybrid filtering perform best for small-, large-, and mid-range queries, respectively, we propose the following heuristic for strategy selection: use Strategy A if $Y\leq\tau_{A}$, Strategy B if $Y\geq\tau_{B}$, and Strategy C if $\tau_{A}<Y<\tau_{B}$. Here, $\tau_{A}$ and $\tau_{B}$ are thresholds distinguishing the optimal ranges for each strategy, derived from historical data analysis. Given these thresholds, we estimate the cardinality of an incoming query to apply the heuristic. Since statistic collection and cardinality estimation are well studied (Harmouch and Naumann, 2017; Oommen and Rueda, 2002) and not the focus of this paper, we provide a simple preprocessing method to validate our heuristic. Specifically, we sample objects from the base dataset as queries and assign each a random query range. We then execute these queries using the three strategies and record recall and latency metrics. Given a recall target, we analyze records meeting this requirement and identify two turning points where pre- and post-filtering outperform hybrid filtering. These points establish $\tau_{A}$ and $\tau_{B}$, guiding strategy selection for future queries.